# LENVIZ: A High-Resolution Low-Exposure Night Vision Benchmark Dataset

### v1

@@ Abstract
Low-light image enhancement is crucial for a myriad of applications, from night vision and surveillance, to autonomous driving. However, due to the inherent limitations that come in hand with capturing images in low-illumination environments, the task of enhancing such scenes still presents a formidable challenge. To advance research in this field, we introduce our Low Exposure Night Vision (LENVIZ) Dataset [^1] Dataset will be released upon publication, a comprehensive multi-exposure benchmark dataset for low-light image enhancement comprising of over 230K frames showcasing 24K real-world indoor and outdoor, with-and-without human, scenes. Captured using 3 different camera sensors, LENVIZ offers a wide range of lighting conditions, noise levels, and scene complexities, making it the largest publicly available up-to 4K resolution benchmark in the field. LENVIZ includes high quality human-generated ground truth, for which each multi-exposure low-light scene has been meticulously curated and edited by expert photographers to ensure optimal image quality. Furthermore, we also conduct a comprehensive analysis of current state-of-the-art low-light image enhancement techniques on our dataset and highlight potential areas of improvement.
Index Terms:
Night Vision, Low Light, Benchmark Dataset, Computer Vision, Computational Photography
I Introduction
@@
Deep learning has emerged as a powerful tool for this task, and while Convolutional Neural Networks (CNNs) can learn complex relationships between low-light and well-lit images, training these models requires large datasets of high-quality image pairs which can be challenging and costly to obtain. Early works [26, 13, 9, 18] used synthetic data with unrealistic lighting conditions, leading to models that struggle with real-world low-light scenarios. Recent efforts[1, 7, 6] focus on creating datasets of real low-light and well-lit image pairs captured under diverse lighting conditions, allowing models to learn generalizable enhancement strategies, but are still limited in size and capture conditions. Having a comprehensive benchmark dataset is essential for advancing research in low-light image enhancement. Such a dataset should provide a diverse range of low-light scenes, capturing various lighting conditions, noise levels, and scene complexities. Moreover, high-quality ground truth images are essential for training and evaluating the performance of different enhancement methods. Unfortunately, existing datasets[26, 1, 7, 6, 18] often suffer from limitations in terms of size, range of capture environments, depicted subjects, or ground-truth availability.
@@
To address these challenges, we introduce the Low Exposure Night Vision (LENVIZ) Dataset, a large-scale multi-exposure benchmark for low-light image enhancement. LENVIZ is the largest to-date (234K frames), up-to 4K resolution, low-light benchmark dataset, offering a comprehensive range of scene types, 3 different imaging sensors, and various ISP settings, addressing a critical gap in existing benchmarks. Additionally, LENVIZ provides 9 distinct exposure frames for each scene along with a long-exposure shot, enabling researchers to explore different enhancement techniques and training schemes. To ensure accurate evaluation and training, we include photography expert-edited ground truth images. Beyond traditional variations, our dataset offers a rich diversity of elements, including different lux levels, capture environments, realistic mannequins for accurate skin tone representation, and facial accessories (glasses, mustaches, etc.), all of these providing more comprehensive features for training (See Fig. 1). To complement the training data, a curated test dataset with 1,468 frames and a total of 203 unique scenes was constructed by leveraging multiple benchmark platforms, including DxOMark [^2] https://corp.dxomark.com/ and internal testing units. This test dataset encompasses a variety of scenes meticulously designed to evaluate not only the performance efficacy of the proposed enhancement and over-exposure recovery methods but also their robustness and generalizability across diverse conditions.
@@
LENVIZ provides a valuable resource for researchers and practitioners working on low-light image enhancement and multi-exposure fusion. By offering a comprehensive, diverse, and high-quality dataset, we enable the development and evaluation of more robust and effective techniques. To demonstrate the value of our benchmark dataset, we conduct an in-depth evaluation of state-of-the-art low-light image enhancement techniques when trained on our dataset. Our evaluation highlights the strengths and weaknesses of existing methods and provides insights into challenges and opportunities for future research in the field. In summary, the contributions of our paper are:
@@
  • Introduction of the LENVIZ Dataset: To the best of our knowledge, we are introducing the largest to-date (234k frame), novel low-light benchmark dataset, offering a diverse range of scene types, imaging sensors, and ISP settings.
@@
  • Flexibility of Inputs: We provide 9 distinct exposure frames for each scene along with a long-exposure shot, enabling researchers to explore different combinations of inputs for their enhancement techniques.
@@
  • Standardized Exposure Calculation: We introduce a meticulous methodology for exposure calculation to determine exposure brackets and lux readings using camera parameters.
  • Curated Test Dataset: We constructed a comprehensive test dataset across 203 unique scenes, leveraging diverse benchmark platforms to evaluate image enhancement and over-exposure recovery methods under various lighting conditions.
  • Benchmarking State-of-the-Art Techniques: We have conducted an in-depth evaluation of current state-of-the-art techniques on the LENVIZ dataset, providing valuable insights into their strengths, weaknesses, and potential areas of improvement.
TABLE I: Summary of existing Low-light enhancement open source datasets. Abbreviations: Indoor/Outdoor: Indoor (I), Outdoor (O), Frame type: Single frame (S), Multi-exposure frames (M), and Long-exposure frame (L)
Dataset Frame Type Resolution Frames Indoor / Outdoor Human subjects Human GT
MIT 5K[26] S 3040x2014 5,000 I,O
NEF[15] S - 88 I,O -
Phos[14] M - 15 I,O
@@ VV Dataset S - 24 I,O
ExDark[6] S - 7,363 I,O
[…]
DarkZurisch[16] S - 5,381 O
ELD[11] M - 60 I
TMDIED[17] M - 222 I
RELLISUR[21] M 0.39-6.25 Mpx 2250 I,O
LDR[10] M - 1800 I,O
UHD-LOL[18] M
@@
LSRW[19] M - 5,650 I,O
Few-shots[4] M+L 512x512 280 I,O
LOM[20] M 3000x4000 25 - 65 I
@@ II Related Work
The increasing demand for AI-driven algorithms to address complex challenges has necessitated the development of diverse, real-world datasets. This is particularly evident in the domains of low-light image enhancement and multi-exposure recovery. At the first stage of low-light imaging, MIT-Adobe FiveK [26] captured 5,000 unpaired RAW images under daylight and low-light conditions in indoor and outdoor settings. In 2013, [15] presented a dataset comprising 88 unpaired low-light images, of which 46 were captured using a Canon digital camera and the remainder were sourced from online websites such as Google, NASA, etc. Although this data set exhibited a notable emphasis on local low-contrast regions with higher global illumination variation, its unpaired nature and limited size hindered its applicability for deep learning training methodologies. Concurrently, [14] introduced the Phos test dataset designed to evaluate feature detectors under varying illumination conditions. Phos included 15 distinct scenes with 9 images captured at uniform illumination levels and 6 images under nonuniform conditions. Despite offering a variety of scenes and multi-exposure intermediate captures, the limited number of images within this paired dataset restricted its utility to testing purposes. Similar data sets have been introduced consisting of a limited number of unpaired frames, namely the VV-dataset [^3] https://sites.google.com/site/vonikakis/datasets, comprising 24 single-exposure unpaired images, and MEF [2] with 17 high-quality multi-exposure images, encompassing underexposed, overexposed, and intermediate illumination levels, and [3] released LIME with low-resolution 10 unpaired low-light data. The small size and unpaired nature of this data limit its use for supervised learning methods.
@@
In 2017, [13] introduced a synthetically generated dataset with over 20,000 images derived from the UCID [22], BSD [23], and Google Image Search datasets. This dataset included 10,000 paired images, each consisting of one high-quality and one low-light natural image. As the first substantial paired low-light benchmark dataset, [13] offered a valuable resource for model training. However, a notable limitation was the absence of camera sensor specifications. [7] then presented the See-In-the-Dark dataset, featuring 4K resolution RAW 12 burst images captured under low-illumination conditions and paired with corresponding long-exposure high-quality images. A total of 424 unique images were acquired using Sony $\alpha$ 7s II and Fujifilm X-T2 cameras. Although the higher diversity and low-light specificity of this dataset were advantageous, its use was restricted by the requirement for RAW input images. [1] introduced the LOw-Light (LOL) dataset, providing 500 paired low-light images. SICE [12] was then introduced with 4,431 multi-exposure paired images using multiple camera modules (6 different DSLR devices $+$ iPhone 6s).
@@
This marked the start of a growing interest in low-light data which led to the introduction of several new low-light datasets. ExDark [6] was released, comprising 7,363 unpaired low-light images captured under twilight conditions. While its size is substantial, its unpaired nature limits its applicability to unsupervised learning. [9] offered a diverse dataset of 10,940 unpaired images, along with 2,500 paired real-world images (1,000 synthetically generated using RAW data from RAISE dataset [24]). [16] contributed 2,920 twilight and 2,461 nighttime images captured using a GoPro Hero 5 camera. Additionally, [11] introduced 60 paired low-light images captured with four different DSLR cameras. Another 222 paired dataset was introduced by [17], providing different lighting conditions (night, sunset, day, cloudy, sunlight etc). [21] introduced a dataset of 2,250 paired images, captured using a Canon EOS 6D camera equipped with a Canon 70-300mm lens. Despite the increasing diversity and realism of newly released datasets, a common limitation was the relatively small size of paired data or the prevalence of large unpaired datasets, which hindered their applicability for most model training paradigms.
@@
In 2023, a significant milestone was reached with the introduction of 4 major paired benchmark datasets offering a substantial number of low-light images. [10] contributed 1,800 unpaired indoor and outdoor RAW images captured with varying ISOs and exposure values, focusing on low-light denoising applications. [18] provided 11,065 paired images. The normal lighting images in this dataset were sourced from publicly available datasets [25], while the corresponding low-light images were synthetically generated. Additionally, LSRW [19], a dataset of 5,650 paired images, was introduced, capturing images with a Nikon D7500 camera and a HUAWEI P40 Pro mobile phone. LSRW consisted of normal/underexposed data pairs, with a fixed ISO of 100 for normal light and 50 for low-light conditions. Finally, [4] released a dataset of 258 paired images captured using three different DSLR cameras.
@@
Despite these advancements, existing datasets often lack a comprehensive combination of multi-exposure capabilities, paired with long-exposure and high-quality ground-truth data, diverse scene locations, lighting ranges, scene types, scene subject types, and camera modules, particularly at up-to 4K resolution. From Table I, we can see that SID[7], ExDark[6], VE-LOL[9], & UHD-LOL[18] were closely able to check all the main categories for the data diversity but still miss either in input frames, size of the dataset, or human-expert editing specific for fixing the image-quality. To address these shortcomings, we analyzed the characteristics and gaps in existing low-light datasets and developed our novel LENVIZ benchmark dataset, incorporating these insights.
III LENVIZ Capture
TABLE II: Summary of the LENVIZ Training Dataset.
Sensor Resolution Human GT Long exposure GT

### v2

@@ Abstract
Low-light image enhancement is crucial for a myriad of applications, from night vision and surveillance, to autonomous driving. However, due to the inherent limitations that come in hand with capturing images in low-illumination environments, the task of enhancing such scenes still presents a formidable challenge. To advance research in this field, we introduce our Low Exposure Night Vision (LENVIZ) Dataset, a comprehensive multi-exposure benchmark dataset for low-light image enhancement comprising of over 230K frames showcasing 24K real-world indoor and outdoor, with-and-without human, scenes. Captured using 3 different camera modules, LENVIZ offers a wide range of lighting conditions, noise levels, and scene complexities, making it the largest publicly available up-to 4K resolution benchmark in the field. LENVIZ includes high quality human-generated ground truth, for which each multi-exposure low-light scene has been meticulously curated and edited by expert photographers to ensure optimal image quality. Furthermore, we also conduct a comprehensive analysis of current state-of-the-art low-light image enhancement techniques on our dataset and highlight potential areas of improvement.
1 Introduction
@@
Deep learning has emerged as a powerful tool for this task, and while Convolutional Neural Networks (CNNs) can learn complex relationships between low-light and well-lit images, training these models requires large datasets of high-quality image pairs which can be challenging and costly to obtain. Early works [MIT5K, LLIE, VE-LOL, UHDLOL2023] used synthetic data with unrealistic lighting conditions, leading to models that struggle with real-world low-light scenarios. Recent efforts [LOL, SID, ExDarkDataset] focus on creating datasets of real low-light and well-lit image pairs captured under diverse lighting conditions, allowing models to learn generalizable enhancement strategies, but are still limited in size and capture conditions. Having a comprehensive benchmark dataset is essential for advancing research in low-light image enhancement. Such a dataset should provide a diverse range of low-light scenes, capturing various lighting conditions, noise levels, and scene complexities. Moreover, high-quality ground truth images are essential for training and evaluating the performance of different enhancement methods. Unfortunately, existing datasets [MIT5K, LOL, SID, ExDarkDataset, UHDLOL2023] often suffer from limitations in terms of size, range of capture environments, depicted subjects, or ground-truth availability.
@@
To address these challenges, we introduce the Low Exposure Night Vision (LENVIZ) Dataset, a large-scale multi-exposure benchmark for low-light image enhancement. LENVIZ is the largest low-light benchmark datasdet to-date (234K frames), offering a comprehensive range of scene types, 3 different camera modules, and various ISP settings, addressing a critical gap in existing benchmarks. Additionally, LENVIZ provides 9 distinct exposure frames for each scene along with a long-exposure shot, enabling researchers to explore different enhancement techniques and training schemes. To ensure accurate evaluation & training, we include photography expert-edited ground truth images. Beyond traditional variations, our dataset offers a rich diversity of elements, including different illuminance levels, capture environments, realistic mannequins for accurate skin tone representation, and facial accessories (glasses, mustaches, etc.), all of these providing more comprehensive features for training (See Fig. LABEL:fig:teaser). We also introduce a curated test dataset with 1,468 frames and a total of 203 unique scenes designed by taking into account multiple benchmark platforms, including DxOMark [^1] https://corp.dxomark.com/. This test dataset encompasses a variety of scenes designed to evaluate not only the performance efficacy of state-of-the-art enhancement and over-exposure recovery methods but also their robustness and generalizability across diverse conditions.
@@
LENVIZ provides a valuable resource for researchers and practitioners working on low-light image enhancement and multi-exposure fusion. By offering a comprehensive, diverse, and high-quality dataset, we enable the development and evaluation of more robust and effective techniques. To demonstrate the value of our benchmark dataset, we conduct an in-depth evaluation of state-of-the-art low-light image enhancement techniques when trained on our data. Our evaluation highlights the strengths and weaknesses of existing methods and provides insights into challenges and opportunities for future research in the field. In summary, the contributions of our paper are:
@@
  • Introduction of the LENVIZ Dataset: To the best of our knowledge, we are introducing the largest to-date (234k frame), novel low-light benchmark dataset.
@@
  • Flexibility of Inputs: We cater 9 distinct exposure frames per scene along with a long-exposure shot, allowing for flexible input combinations.
@@
  • Standardized Exposure Calculation: We introduce a method to determine exposure brackets and illuminance readings using camera parameters.
  • Curated Test Dataset: We constructed a comprehensive test dataset across 203 unique scenes, leveraging diverse benchmark platforms to evaluate enhancement techniques under varying conditions.
  • Benchmarking State-of-the-Art Techniques: We have conducted an in-depth evaluation of current state-of-the-art techniques in the LENVIZ dataset.
Table 1: Summary of existing Low-light enhancement open source datasets. Abbreviations: Indoor/Outdoor: Indoor (I), Outdoor (O), Frame type: Single frame (S), Multi-exposure frames (M), and Long-exposure frame (L)
Dataset Frame Type Resolution Frames Environment Human subjects Human GT
MIT 5K[MIT5K] S 3040x2014 5,000 I,O
NEF[NEF] S - 88 I,O -
Phos[Phos] M - 15 I,O
@@ VV Dataset S - 24 I,O
ExDark[ExDarkDataset] S - 7,363 I,O
VE-LOL[VE-LOL] M $\|$  S 1080x720 2,500 $\|$  10,940 I,O
DarkZurisch[DarkZurisch] S - 5,381 O
ELD[ELD] M - 60 I
TMDIED[TMDIED2021] M - 222 I
RELLISUR[rellisur2023] M 0.39-6.25 Mpx 2250 I,O
LDR[LDR] M - 1800 I,O
UHD-LOL[UHDLOL2023] M
@@
LSRW[LSRW2023] M - 5,650 I,O
Few-shots[NikonfewShot] M+L 512x512 280 I,O
LOM[LOM2023] M 3000x4000 25 - 65 I
SDE[eventImage] M 346x260 30,000 I,O
RLED[NERNet] M - 80,400 I,O
@@ 2 Related Work
The increasing demand for AI-driven solutions to address complex challenges has spurred the development of diverse, real-world datasets. This is particularly evident in the domains of low-light image enhancement and multi-exposure recovery. MIT-Adobe FiveK [MIT5K] captured 5,000 unpaired RAW images under daylight and low-light conditions in indoor and outdoor settings. [NEF] presented a dataset comprising 88 unpaired low-light images, of which 46 were captured using a Canon digital camera and the remainder were sourced from online websites such as Google, NASA, etc. Although this dataset exhibited a notable emphasis on local low-contrast regions with higher global illumination variation, its unpaired nature and limited size hindered its applicability for deep learning training methodologies. The Phos[Phos] dataset provides 15 distinct scenes with 9 images captured at uniform illumination levels and 6 images under nonuniform conditions. Despite offering diverse and extensive scenes and multi-exposure intermediate captures, the limited number of images within this paired dataset restricted its utility to testing purposes. Similar-sized datasets include the VV-dataset [^2] https://sites.google.com/site/vonikakis/datasets with 24 single-exposure unpaired images, and MEF [MEF] with 17 high-quality multi-exposure images, encompassing underexposed, overexposed, and intermediate illumination levels, and LIME [LIME] with low-resolution 10 unpaired low-light data.
@@
LLIE [LLIE] introduced a synthetically generated dataset with more than 20,000 images derived from the UCID [ucid], BSD [BSD], and Google Image Search datasets. This included 10,000 paired images, each consisting of one high-quality and one low-light natural image. As the first substantial paired low-light benchmark dataset, it offered a valuable resource for model training; however, a notable limitation was the absence of camera module specifications. The See-In-the-Dark dataset [SID], features RAW 12 burst images captured under low-illumination conditions and paired with corresponding long-exposure high-quality images. A total of 424 unique images were acquired using Sony $\alpha$ 7s II and Fujifilm X-T2 cameras. Although the higher diversity and low-light specificity of this dataset were advantageous, its use was restricted by the requirement for RAW input images. The LOw-Light (LOL) dataset [LOL], provides 500 paired low-light images, and SICE [SICE] then introduced 4,431 multi-exposure paired images using multiple camera modules (6 different DSLR devices ++ iPhone 6s).
@@
This marked the start of a growing interest in low-light data which led to the introduction of several new low-light datasets. ExDark [ExDarkDataset] was released, comprising 7,363 unpaired low-light images captured under twilight conditions. Although its size is substantial, its unpaired nature limits its applicability to unsupervised learning. [VE-LOL] offered 10,940 unpaired images, along with 2,500 paired real-world images (1,000 synthetically generated using RAW data from RAISE dataset [raise]). [DarkZurisch] contributed 2,920 twilight and 2,461 night-time images captured using a GoPro Hero 5 camera. Additionally, [ELD] introduced 60 paired low-light images captured with four different DSLR cameras. Another 222 paired dataset was introduced by [TMDIED2021], providing different lighting conditions (night, sunset, day, cloudy, sunlight etc). [rellisur2023] introduced a dataset of 2,250 paired images, captured using a Canon EOS 6D camera equipped with a Canon 70-300mm lens. Despite the increasing diversity and realism of newly released datasets, a common limitation was the relatively small size of paired data or the prevalence of large unpaired datasets, which hindered their applicability for most model training paradigms.
@@
In 2023, a significant milestone was reached with the introduction of 4 major paired benchmark datasets offering a substantial number of low-light images. [LDR] contributed 1,800 unpaired indoor and outdoor RAW images captured with varying ISOs and exposure values, focusing on low-light denoising applications. [UHDLOL2023] provided 11,065 paired images, sourcing the normal lighting images from publicly available datasets [publicData] and synthetically generating the corresponding low-light images. LSRW’s[LSRW2023] 5,650 paired images were also released, the images captured with a Nikon D7500 camera and a HUAWEI P40 Pro phone. LSRW consisted of normal/underexposed data pairs, with a fixed ISO of 100 for normal light and 50 for low-light conditions. In the most recent releases, author working on event-based method such as [eventImage] provided SDE dataset that comprises of 30K paired indoor and outdoor images with dark and normal illumination levels. Additionally,  [NERNet] released some large-scale (80K), multi-illumination levels and pixel-aligned GT for low-light conditions.
@@
Despite these advancements, existing datasets often lack a comprehensive combination of multi-exposure capabilities, paired with long-exposure and high-quality ground-truth data, diverse scene locations, lighting ranges, scene types, scene subject types, and camera modules, particularly at up-to 4K resolution. From Table 1, we can see that while some datasets provide diverse data, they are either lacking in multi-frames inputs, dataset size, or human-generated ground-truth. To address these shortcomings, we analyzed the characteristics and gaps in existing low-light datasets and developed our novel LENVIZ benchmark dataset, incorporating these insights.
3 LENVIZ Capture
Our dataset is composed of $24K$ low-light scenes captured over $3$ different camera modules. We aimed to create a diverse dataset suitable for pixel-level image processing tasks such as image enhancement and high dynamic range (HDR) imaging in challenging low-light conditions. With this in mind, we provide up to 9 multi-exposure frames as well as a long-exposure shot for each scene ($234,688$ total frames) to aid in the development of single, and multi-exposure image enhancement approaches. In this section we detail the capture pipeline for our data as well as camera-specific details and configurations. Additional details on the relationship between the estimated illuminance and the exposure time are provided in Supp 7.1.2
3.1 Camera module Information
To broaden our dataset’s representativeness and improve the generalizability of algorithms trained using our data, we provide scenes captured over $3$ different camera modules. Information on the specific details (aperture, camera module-size, FOV, etc.) is provided in Supp 7.1. While our S5K4H7YX03-FGX9 camera is better suited for close range captures, S5KJN1SQ03 and S5KJNS camera features allow for richer capture of details at medium and long range.
It is important to note that, the performance and characteristics of camera module can vary based on their placement on a device and intended function, making certain camera module more suitable for specific applications. Our first camera module, S5K4H7YX03-FGX9, was the front-facing camera on a mobile phone. These cameras are essential for selfies and video calls but are constrained in physical size due to their placement above the phone display. Although their resolution is generally lower than that of rear cameras, they are designed to capture scenes at close range (usually an arm’s length) and often benefit from additional light from the screen. As such, our images from this camera module tend to contain a higher concentration of faces (distribution presentation in Supp 7.2.1). The other two camera module, S5KJN1SQ03 and S5KJNS, were used as rear cameras on different devices. These are designed to serve as the primary lens for capturing the environment. Unlike front cameras, they offer higher resolution and specifications for more detailed scenes, which typically involve greater depth and lower lighting conditions. While face detection remains important for rear cameras, they are also used for capturing a wider range of subjects, such as landscapes and objects.
3.2 Ambient Light Illuminance Calculation
For our data acquisition, we implemented stringent controls over camera parameters such as ISO sensitivity, exposure time, and aperture. Concurrently, we performed an estimation of the scene illuminance levels using our capturing devices, this measurement indicates the intensity of light in a scene and is essential for characterizing low-light environments and is included for each scene. This approach ensured a highly controlled environment for image capture, thereby enhancing the dataset’s suitability for benchmarking low-light image enhancement algorithms. We estimate the illuminance readings of the captured images using key camera parameters like exposure time, ISO sensitivity, and lens aperture:
$\text{L'}=\beta_{\text{illuminance}}\times f^{2}\times\left(\frac{10^{11}}{t_{\text{exp}}\times ISO}\right)$ (1)
Here, the exposure time ($t_{\text{exp}}$) determines the duration (in $ns$) that the camera module is exposed to light. $ISO$ sensitivity controls the camera module’s sensitivity to light, with higher values increasing sensitivity but also introducing noise. Lens aperture ($f$) regulates the amount of light reaching the camera module (fixed for all camera modules in our dataset). Finally, camera-specific tuning constants ($\beta_{\text{illuminance}}$) account for camera module characteristics and ISP tuning. Our illuminance estimation is inversely proportional to both the exposure time and the ISO value. Longer exposure times and higher ISO values reduce the calculated illuminance. The aperture size $f$, directly influences the light-gathering capacity of the lens. Additionally, when a post-processing sensitivity boost is applied (e.g., through ISO adjustment as part of the ISP tuning of a camera module), the ISO value is modified by the post-boost percentage.
3.3 Long-Exposure Time Calculation
The exposure time for our long-exposure images is computed based on the ambient light intensity, measured in illuminance, using a formula derived from the relationship between illuminance and exposure time. Equation 2 governs the exposure time calculation for the long-exposure ground truth image given the estimated illuminance of the scene as follows:
$\text{exposure\textsubscript{gt}}=\begin{cases}\frac{\gamma_{0}}{\text{lux}}&\text{if }\text{illuminance}<1\\ \frac{\gamma_{1}}{\text{lux}}&\text{if }1<\text{illuminance}<8\\ \frac{\gamma_{2}}{\text{lux}}&\text{if }8\leq\text{illuminance}<10\\ \frac{\gamma_{3}}{\text{lux}}&\text{if }10\leq\text{illuminance}<15\\ \frac{\gamma_{4}}{\text{lux}}&\text{if }\text{illuminance}\geq 15\end{cases}$ (2)
The $\gamma$ values of the long-exposure settings is determined empirically and the values we used during our collection are $\gamma_{1}=48,\gamma_{2}=48,\gamma_{3}=52,\gamma_{4}=60$ . The detailed explanation on the process and experimental setup we used to determine the optimal $\gamma$ values are provided in Supp 7.1.1.
Table 2: Summary of the LENVIZ Training Dataset.
Camera Module Resolution Human GT Long exposure GT